OpenStudy (anonymous):

simplify again: x^2+4x+4/x^2-4

OpenStudy (a_clan):

(x+2)/(x-2)

OpenStudy (anonymous):

x^2+4x+4/x^2-4 =((x+2)/(x+2))/(x+2)/(x-2)) then you cancel out the numerator and denominator for (X+2) =(x+2)/(x-2) fan me if you need any more algebra help

OpenStudy (anonymous):

x^2+4x+4/x^2-4=(x+2)(x+2)/(x+2)(x-2) =(x+2)/(x-2)